# Robustness of LiDAR-Based Pose Estimation: Evaluating and Improving Odometry and Localization Under Common Point Cloud Corruptions

### v2

@@ IV Experiment Setup
IV-A Subject LiDAR-Based Pose Estimation Systems
TABLE IV: Subject Systems. The errors are reported based on our replication.
Category Dataset Year Method Error
Odometry MULLS [24] 2021 Feature-based 0.56%
KISS-ICP [15] 2023 Direct 0.51%
Delora [19] 2021 Learning-based 7.78%
NeRF-LOAM [22] 2023 NeRF-based 1.70%
Localization LocNDF [21] 2023 NeRF-based 0.059m
We experiment with four SOTA LiDAR odometry and one LiDAR localization systems, as shown in Table IV. We cover one representative work from different types of mainstream point cloud registration methods, i.e., feature-based, direct, and learning-based methods [10]. We also include two neural distance field-based approaches, which have become popular since the success of the NeRF [20].
MULLS [24] is a feature-based LiDAR odometry that uses ground filtering and principal components analysis to extract features from the point cloud. It builds local sub-maps using the previous point cloud and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose. KISS-ICP [15] is a direct matching LiDAR odometry that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling. Delora [19] is a learning-based, unsupervised LiDAR odometry. It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image. Then it predicts the translation and rotation by feeding the features into a neural network. NeRF-LOAM [22] employs a neural network to predict the signed distance field (SDF). It first builds the previous frames into an octree. Then it samples intersected points between laser rays and voxels in the previously built octree. Given sampled points, the neural network predicts the SDF. Since sampled points are transformed to the same coordinate system as the previously built octree by the pose to be estimated, NeRF-LOAM optimizes the pose, neural network, and voxel embedding jointly by minimizing the loss between predicted SDF and real SDF. Different from NeRF-LOAM, LocNDF trains a neural network offline to predict the SDF. Given the online LiDAR scan, it tries to find a pose $P$ so that the neural network can produce minimal SDF when sampled points are transformed by $P$.
@@ IV-B Datasets
We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [41] to evaluate the four odometry systems, Delora, MULLS, KISS-ICP, and NeRF-LOAM. KITTI 00 to 10 sequences are released with ground truths. We conducted our experiment on sequences 09 and 10 because (1) Delora was trained on sequences 00-08 hence only 09 and 10 remain. (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
Apollo South-Bay [26] contains completed ADS recordings covering various driving scenarios. In this work, we used a subset of ColumbiaPark-3, consisting of 700 testing frames, which is consistent with LocNDF’s pre-trained model [21].
@@ IV-C Evaluation Metrics
For the evaluation of the precision of pose estimation, we adopted the relative pose error (RPE) [41]. RPE is widely used in previous work [10] to measure and compare the performance of pose estimation. RPE is calculated as:
$RPE_{rot}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}\angle[(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})]$ (2)
$RPE_{trans}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}||(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})||_{2}$ (3)
where $RPE_{trans}$ measures the relative translation error and $RPE_{rot}$ measures the relative rotation error between the estimated pose and the ground truth, $\mathscr{F}$ is a set of frames $(i,j)$, $\hat{p}$ is the estimated pose and $p$ is the ground truth pose, $\ominus$ is the inverse compositional operator [42] and []delimited-[$\angle$[$\cdot$]∠ [ ⋅ ] is the rotation angle.
@@ V Experiment Results
V-A Evaluating Robustness of LiDAR-Based Pose Estimation
@@
Method. We applied our robustness framework to evaluate the subject odometry systems. We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I. Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish one a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
Figure 2: Performance of each system on corrupted KITTI LiDAR Odometry Sequence 09 and 10. Errors with significantly different magnitudes from the majority are aligned at their magnitude and the actual values are labeled in the figure. For example, an error of 12.79% is aligned to “$>$10.0”. If multiple errors overlap at the same magnitude, we label the two highest severity levels.
Results. We show the results of RPEtrans in Figures 2. Due to the space limitation, we do not present the RPErot in the paper as it is correlated with RPEtrans. We share the detailed results of RPErot in our public artifact.
For the non-learning systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data). Furthermore, each system is vulnerable to a particular type of perturbation, where their performance drops significantly. The error of MULLS increases from below 0.5% to up to 0.8% on most of the corruptions. However, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10. Similarly, kiss-icp is especially vulnerable to “background noise”. While the error only increased by less than 0.2% in most cases, the error of kiss-icp caused by background noise increases up to 80% (with severity 5 on KITTI 10) from less than 0.6% RPEtrans. Moreover, even the lightest background noise (severity 1) causes a significant error increase from less than 0.6% to more than 3% RPEtrans on KITTI 10.
We observe different results for the learning-based systems, Delora and NeRF-LOAM. Delora is vulnerable to most perturbations, especially those related to noise. The RPEtrans error increases from less than 10 percent to tens of percent (up to 50%). Lastly, NeRF-LOAM is quite different from others. We find the results differ very much between the two KITTI sequences, indicating that NeRF-LOAM’s robustness results highly depend on data. In KITTI 09, NeRF-LOAM is not sensitive to most perturbations except for one. While in KITTI 10, NeRF-LOAM is shown to be highly sensitive in most cases. Also, interestingly, in many cases, it shows better robustness with a higher severity than a lower severity. For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well with severity 5 which is close to the baseline. NeRF-LOAM optimizes a mini-network online for tens of iterations. We hypothesize that the poor and unstable results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
For the LiDAR localization system LocNDF, our evaluation shows that it is very robust against all types of corruption. The robustness is at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
Our results show that the SOTA LiDAR localization system, LocNDF, is robust against all point cloud corruptions. SOTA LiDAR odometry systems are vulnerable to one or multiple types of point cloud corruptions.
V-B Improving Robustness of LiDAR-Based Pose Estimation
Method. We only concern MULLS, kiss-icp, and Delora in this denoising experiment, since LocNDF is robust against all kinds of corruption, while NeRF-LOAM could not produce consistent results between different trials on the same data. Firstly, we apply the bilateral filter to denoise the corrupted data and then re-run all the algorithms on the denoised data. Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying the proposed point cloud corruptions and re-train Delora on the augmented data.
We focus on denoising the most representative corruption types that affect the algorithms most, i.e., local density decrease (MULLS and Delora), background noise (kiss-icp), uniform noise (Delora), Gaussian noise (Delora), and rain/snow (moderate effects on all systems). We denoise these types of corruption at severity 1, 3, and 5. We also denoise the clean data to see if the denoising has side effects on clean data.
@@ V-B Improving Robustness of LiDAR-Based Pose Estimation
Similarly, we fine-tune Delora on the corruptions that affect it most significantly, namely, rain, snow, fog, upsample, uniform noise, Gaussian noise, impulse noise, local density decrease, cutout, and layer deletion. We only augmented the training data with these corruptions at severity 5 and fine-tuned Delora for 6 epochs with the default training hyper-parameters.
Results.
Figure 3: (a) Performance on the denoised data. (b) Performance of the retrained model.
We present the performance of the three systems on the denoised data for KITTI 09 in Figure 3 (a). The results on KITTI 10 are similar and available in our replication package. For MULLS, local density decrease (local_dec) at severity 5 is the corruption that affects it most significantly. We find that the errors even increased by 2.43% after denoising. Besides, we find that denoising could reduce the error by 0.26% for Gaussian noise. The changes are insignificant for the remaining cases (less than 0.1%).
For kiss-icp, background noise (bg_noise) is the corruption that affects it most significantly. We find that the denoising can reduce the error by 0.2% (severity 1) to 3.22% (severity 5), that is, the negative effect of the corruption is fully eliminated (we observed the same effects on KITTI 10, where the errors at severity 5 are reduced by 80%). However, the denoising causes the errors on rain and snow to increase by 0.13% to 0.18%, and on local density decrease at severity 5 to increase by 0.28%. The changes are insignificant for the remaining cases.
For Delora, we find that denoising can reduce its errors in most cases including the baseline, by 0.44% to 16.08%. Similar to MULLS and kiss-icp, we find that the errors in local density decrease (local_dec) increased by 2.74%.
Bilateral filter can effectively mitigate noise-induced corruption in general, and fully eliminate the negative effects of background noise. However, it does not help with density-related corruption (local density decrease).
@@ Results.
We present the performance of the Delora after fine-turning for KITTI 09 in Figure 3 (b). We find that re-training can bring a much higher improvement than denoising: the errors are reduced by 2.73% to 35.79% for the baseline and all types of corruption. However, for KITTI 10, despite the overall effects being similar, we observed a slight increase in the errors for the baseline and a few corruptions. It is important to note that our fine-tuning was relatively naive: we only augmented KITTI 00 to 08 with the selected corruptions at severity 5 and used the original learning rate. We believe that further improvements could be achieved by a more refined fine-tuning process, such as augmenting the dataset with the corruption at different severities, and adjusting the learning rate and the ratio between clean and corrupted data.
@@
Re-training with the corruption data is a more effective way to improve the robustness of the learning-based LiDAR pose estimation system (Delora).
@@
In this work, we extensively evaluated the robustness of LiDAR-based pose estimation systems, including odometry and localization, against 18 synthetic real-world point cloud corruptions. Our findings indicate that all SOTA LiDAR-based odometry systems are particularly vulnerable to certain corruption types. We further show that simple bilateral filtering effectively mitigates noise-induced data corruptions, however, density-related corruptions remain challenging. Finally, we show that corruption-based data augmentation not only significantly enhances the robustness of learning-based LiDAR odometry systems but also improves their performance on clean data.

### v3

@@ IV Experiment Setup
IV-A Subject LiDAR Pose Estimation Systems
Category Name Year Method
Odometry MULLS 2021 Feature-based
KISS-ICP 2023 Direct
DeLORA 2021 Learning-based
NeRF-LOAM 2023 Learning-based
Localization LocNDF 2023 Learning-based
TABLE IV: Subject LiDAR odometry and localization techniques.
We evaluate four odometry and one localization systems, as shown in Table IV. The subject systems comprise two representative classical methods, MULLS and kiss-icp, and three learning-based approaches, DeLORA, NeRF-LOAM, and LocNDF. MULLS [20] is a feature-based LiDAR odometry that uses ground filtering and principal components analysis to extract features from the point cloud. It builds local sub-maps using the previous point cloud and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose. KISS-ICP [4] is a direct matching LiDAR odometry that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling. DeLORA [7] is a learning-based, unsupervised LiDAR odometry. It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image. Then it predicts the translation and rotation by feeding the features into a neural network. NeRF-LOAM [18] employs a neural network to predict the signed distance field (SDF). It first builds the previous frames into an octree. Then it samples intersected points between laser rays and voxels in the previously built octree. Given sampled points, the neural network predicts the SDF. Since sampled points are transformed to the same coordinate system as the previously built octree by the pose to be estimated, NeRF-LOAM optimizes the pose, neural network, and voxel embedding jointly by minimizing the loss between predicted SDF and real SDF. Different from NeRF-LOAM, LocNDF trains a neural network offline to predict the SDF. Given the online LiDAR scan, it tries to find a pose $P$ so that the neural network can produce minimal SDF when sampled points are transformed by $P$.
@@ IV-B Datasets
We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [41] to evaluate the four odometry systems, DeLORA, MULLS, KISS-ICP, and NeRF-LOAM. The reason is that KITTI is one of the most popular dataset for odometry and all of the selected odometry systems have been evaluated on KITTI. KITTI 00 to 10 sequences are released with ground truths. We conducted our experiment on sequences 09 and 10 because (1) DeLORA was trained on sequences 00-08 hence only 09 and 10 remain. (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
We use Apollo South-Bay [24] to evaluate LocNDF. Apollo South-Bay contains completed ADS recordings covering various driving scenarios. LocNDF released the pretrained model and pre-built map on this dataset. In this work, we used a subset of ColumbiaPark-3, consisting of 700 testing frames, which is consistent with LocNDF’s pre-trained model [17].
@@ IV-C Evaluation Metrics
For the evaluation of the precision of pose estimation, we adopted the translational relative pose error (RPEtrans[41]. RPEtrans is widely used in previous work [13] to measure and compare the performance of pose estimation. RPEtrans is calculated as:
[…]
where $\mathscr{F}$ is a set of frames $(i,j)$, $\hat{p}$ is the estimated pose and $p$ is the ground truth pose, $\ominus$ is the inverse compositional operator [42]. A pose $p$ consists of a rotation matrix $R$ and a translation vector $t$.
@@ V Experiment Results
V-A RobustLOL: Robustness Results of LiDAR Odometry and Localization
@@
Method. We applied our robustness framework to evaluate the subject odometry and localization systems. We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I. Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24 GB GPU memory and several days to finish a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
Figure 3: Performance of each system on KITTI 09.
Figure 4: Performance of each system on KITTI 10.
Results. Figure 3 and 4 present RPEtrans on KITTI 09 and 10 for all corruptions of severities 1, 3, and 5. We omit the results of LocNDF as we observe no performance degradation under any corruption in our experiment. We share the complete results in the released artifacts.
Non-learning systems show overall robustness to most perturbations with each system particularly sensitive to certain corruption types, leading to significant performance drops. MULLS is sensitive to local density decrease (local dec) at severity 5 where RPEtrans increases from 0.3% to over 18% and kiss-icp is sensitive to background noise where RPEtrans rises sharply from less than 0.6% on the clean data to over 80%. Our investigation shows that local density reduction disrupts key point clusters which MULLS relies on, making its feature extraction unreliable. On the other hand, background noise introduces uniformly distributed noisy points which disrupt kiss-icp’s direct point-on-point registration.
In contrast, learning-based odometry systems, i.e., DeLORA and NeRF-LOAM are more sensitive to most corruptions. Notable exceptions include background noise, spherical-coordinate noises (uniform, Gaussian and impulse noises (SCS)), local density increase, and beam deletion, where performance remains relatively stable due to DeLORA’s 2D range image representation. During projection, multiple 3D points may map to the same pixel, with only the closest point retained, which significantly limit the effect of these perturbations. Thus, background noise appears as mild pixel-level noise, spherical-coordinate noise alters depth without affecting pixel locations, beam deletion has limited impact due to point cloud redundancy, and local density increase adds points that often overlap with existing pixels, preserving image structure. In contrast, other corruptions introduce geometric distortions or remove key features, leading to more severe performance drops. Finally, we find that NeRF-LOAM shows substantial instability, particularly on KITTI 10. In some cases, increasing corruption severity produces lower errors which is counterintuitive. To investigate this, we repeated the experiments under each corruption two additional times (three runs in total) with identical inputs. The resulting trajectories and final errors vary substantially across runs. We attribute this variability to random network initialization and the limited number of online optimization iterations.
Our systematic evaluation reveals key vulnerabilities in various LiDAR odometry systems.
V-B Detection-and-Filter Pipeline
Figure 5: (a) $RPE_{trans}$ with our detection-and-filter pipeline. (b) $RPE_{trans}$ of the fine-tuned DeLORA on the augmented dataset. Colored bars (light gray, medium gray, dark gray) report the RPE at severity levels 1, 3, and 5 after applying our pipeline or fine-tuning; the hatched white bars with “/” show the corresponding errors without our pipeline or fine-tuning.
Method. We apply the detection-and-filter pipeline on all systems except NeRF-LOAM because it could not produce consistent results between different trials on the exact same data and LocNDF because it is robust against all proposed corruptions.
We implement the classifier proposed in Section III-B1 using PyTorch 1.13.1+cu117 and train it to recognize eight representative corruptions: rain, snow, fog, uniform noise (CCS), Gaussian noise (CCS), background noise, local density increase, and local density decrease. Due to their similar scattering effects, we merge rain and snow into one class; likewise, uniform and Gaussian noise (CCS) are grouped together for their comparable distortion patterns. We train the model on KITTI 09 (80% train, 20% validation) for 10 epochs using a batch size of 8, a learning rate of 0.001 with the Adam optimizer. We apply it on KITTI 10 and we use the bilateral filter to restore the noise-corrupted PCD. All experiments are run on Ubuntu 22.04 LTS with an AMD Ryzen 32-core CPU, 256 GB RAM, and an NVIDIA GeForce RTX 3090 GPU.
Results. Our classification model achieves an average accuracy of 95.54%, as shown in Table V. We notice that clean and fog have a comparably lower accuracy because they are often misclassified as local density increase. The reason is that (i) fog has sparse reflections which increase the points density, and (ii) occasional high-reflecting surfaces in clean data also increases points density.
Weather Noise Density
Clean Fog R+S U+G Back Inc Dec
Acc 90.02% 80.83% 98.32% 100% 100% 100% 97.84%
TABLE V: Classification results for seven class of noises, i.e., Clean, Fog, Rain+Snow (R+S), Uniform+Gaussian Noise (U+G), Background Noise (Back), Local Density Increase (Inc), and Local Density Decrease (Dec)
We present the RPEtrans of the three systems incorporated with our pipeline on KITTI 10 in Figure 5 (a). MULLS is shown to be robust against most noise in the previous section. Applying our detection-and-filter pipeline to the noisy data causes negligible performance degradation (less than 0.1%). kiss-icp significantly benefits from the detection-and-filter pipeline, with post-denoising errors returning to the clean-data baseline. DeLORA also has a significant error reduction after filtering noises in data, however, the errors are still much higher than the baseline.
Our classifier could classify noise types accurately and our detection-and-filter pipeline could significantly reduce the error on noise-corrupted point cloud data.
V-C Fine-Tuning
@@
Method. We fine-tune the learning-based system, i.e., DeLORA, on the dataset augmented with the corruptions that affect it most significantly, namely, rain, snow, fog, upsample, uniform noise (CCS), Gaussian noise (CCS), impulse noise (CCS), local density decrease, cutout, and layer deletion. We exclude NeRF-LOAM and LocNDF for the same reasons mentioned in Section V-B.
@@
Results. We present the performance of the fine-tuned DeLORA on corrupted KITTI 10 scans (KITTI 09 results are available in the replication package) in Figure 5 (b). We find that the errors are reduced by up to 35.79% for all types of corruption after fine-tuning which brings the performance close to the clean data. On the baseline, the fine-tuned DeLORA achieves a 2.7% decrease in RPEtrans on KITTI 09 but a 3.3% increase on KITTI 10 compared to the original model. It is important to note that our fine-tuning was comparably simple: we only augmented KITTI 00 to 08 with the selected corruptions at severity 5 and used the original learning rate. We believe that further improvements could be achieved by a more refined fine-tuning process, such as augmenting the dataset with the corruption at different severities, and adjusting the learning rate and the ratio between clean and corrupted data.
@@
Fine-tuning with the corruption data is an effective way to improve the robustness of the learning-based LiDAR odometry system (DeLORA).
@@ VI Conclusion
In this work, we propose RobustLOL, a systematic framework for evaluating the robustness of LiDAR-based odometry and localization systems against 18 types of synthetic real-world point cloud corruptions. Our evaluation reveals critical vulnerabilities in different types of odometry systems, with performance degradation of up to 80%. We further demonstrate that our simple detection-and-filter pipeline can effectively mitigate noise-induced corruptions, restoring performance to near-clean levels. Finally, we show that corruption-based data augmentation substantially enhances the robustness of learning-based systems against point cloud corruptions. Notably, it also improved clean-data performance on KITTI 09, though a slight decrease was observed on KITTI 10.